OpenStudy (anonymous):

Which of the following represents the graph of f(x) = 2x + 2? i will give a medal if you help cooperatively

OpenStudy (anonymous):

you dont have a graph? using the y=mx+b formula m=2 and b=2

OpenStudy (anonymous):

here you go

OpenStudy (whpalmer4):

None of those are f(x) = 2x+2! f(x) = 2x + 2 is a straight line.

OpenStudy (anonymous):

I got the same thing that's why I had to question it

OpenStudy (whpalmer4):

well, look carefully at the equation, it's 2^x + 2, not 2x + 2

OpenStudy (whpalmer4):

so let's compute a few values of f(x) and see which graph has them all on the curve. if x = 0, f(x) = f(0) = 2^0 + 2 =

OpenStudy (anonymous):

how would I graph this?

OpenStudy (anonymous):

its the last one

OpenStudy (anonymous):

x^2 look at it again thats why i couldn't fiqure it out

OpenStudy (anonymous):

how do you know because I dont want guessing this is critical

OpenStudy (whpalmer4):

let's work through this. I asked you a question: if x=0, f(x) = f(0) = 2^0 + 2 = what is the number that goes after the final equals sign?

OpenStudy (whpalmer4):

@binks it isn't x^2, it's 2^x

OpenStudy (anonymous):

f(x) = 2^x + 2 typo sorry

OpenStudy (anonymous):

f(x) = 2^x + 2

OpenStudy (whpalmer4):

@jedi3000 follow along with the steps I give you and you'll know exactly how to do it and be 100% confident in your answer.

OpenStudy (anonymous):

alright

OpenStudy (whpalmer4):

so, for the last time, if x=0, f(x) = f(0) = 2^0 + 2 = what is the number that goes after the final equals sign?

OpenStudy (anonymous):

3 goes after the equal sign

OpenStudy (anonymous):

What transformation has changed the parent function f(x) = 3(2)x to its new appearance shown in the graph below?

OpenStudy (anonymous):

that graph I sent you was for the other question sorry

OpenStudy (whpalmer4):

right, 2^0 + 2 = 1 + 2= 3, so at x = 0, y = f(x) = f(0) = 3 now let's find the value of the curve at x = 1 f(x) = f(1) = 2^1 + 2 = and at x = 2: f(x) = f(2) = 2^2 + 2 = there should be only one graph that has those points

OpenStudy (whpalmer4):

f(x) = 3(2)^x you can't just type the symbols you see, you have to add other symbols as necessary to indicate how the symbols are positioned! 3(2)x is NOT the same as 3(2)^x

OpenStudy (whpalmer4):

we go about this pretty much the same way as graphing that other equation. pick a couple of points, x = 0 is almost always a good one, and so is x = 1. find the value of the function at those points. f(x) = f(0) = 3(2)^0 = 3*2^0 = 3*1 = 3 f(x) = f(1) = 3(2)^1 = 3*2^1 = 3*2 = 6 f(x) = f(2) = 3(2)^2 = 3*2^2 = 3*4 = 12

OpenStudy (whpalmer4):

Next, we plot those points and see if we can recognize how the curve has been shifted. Here we have 2 basic transformations to choose from: f(x) + 2 and f(x) + 4 both shift the curve up (subtracting a value from f(x) would shift it down) f(x+2) and f(x+4) both shift the curve sideways. A feature of the curve previously found at x = 0 now happens at x = -2 or x = -4, shifting the curve to the left. f(x-2) or f(x-4) would shift it to the right.

OpenStudy (whpalmer4):

you can also have a combination (though none of the answer choices here are a combination). f(x-1) + 2 shifts the graph of f(x) right 1 unit and up 2 units, for example.

OpenStudy (anonymous):

I got f(x) − 2 for the second problem

OpenStudy (anonymous):

For the first prob the answer I put the last one

OpenStudy (anonymous):

I am not sure if I am correct though.... :-/